5. Given the standard enthalpies of formation for the following substances, determine the change in enthalpy for the combustion of 1.0 mol of propane, C3H8.
Substance ΔHf in kJ/mol
C3H8(g) -103.8
CO2(g) - 393.5
H2O(l) - 285.8
6. Calculate the energy released as heat when 45.8 g of Fe(s) is combined with 27.9g of O2(g) if the ΔH°rxn is -1648 kJ/mol. The balanced equation for the reaction is
4 Fe(s) + 3O2(g) = 2Fe2O3(s)
1
Expert's answer
2019-12-12T08:27:01-0500

5.

Combustion of propane is given by :


"C_3H_8+5O_2\to 3CO_2+4H_2O"

enthalpy of combustion of propane =


"3\times\Delta H_f(CO_2) +4\times \Delta H_f(H_2O)-\Delta H_f(C_3H_8)"

"3\times(-393.5)+ 4\times(-285.8)-(-103.8)\\=2219.2\ kJ/ mol"


6.



"4 Fe(s) + 3O_2 \to 2Fe_2O_3(s)"

Iron is the limiting reagent in this problem

Now,

224 gm of iron releases 1648 kJ energy

then, 45.8 gm of Iron will release = "\frac{45.8}{224}\times1648 =337 \ kJ \ energy"